# An analogue of Reed’s conjecture for digraphs

### v2

@@ Abstract
Reed in 1998 conjectured that every graph $G$ satisfies $\chi(G)\leq\lceil\frac{\Delta(G)+1+\omega(G)}{2}\rceil$. As a partial result, he proved the existence of $\varepsilon>0$ for which every graph $G$ satisfies $\chi(G)\leq\lceil(1-\varepsilon)(\Delta(G)+1)+\varepsilon\omega(G)\rceil$. We propose an analogue conjecture for digraphs. Given a digraph $D$, we denote by $\vec{\chi}(D)$ the dichromatic number of $D$, which is the minimum number of colours needed to partition $D$ into acyclic induced subdigraphs. We let ${\omega}(D)$ denote the size of the largest biclique (a set of vertices inducing a complete digraph) of $D$ and $\tilde{\Delta}(D)=\max_{v\in V(D)}\sqrt{d^{+}(v)\cdot d^{-}(v)}$. We conjecture that every digraph $D$ satisfies $\vec{\chi}(D)\leq\lceil\frac{\tilde{\Delta}(D)+1+\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)}{2}\rceil$, which if true implies Reed’s conjecture. As a partial result, we prove the existence of $\varepsilon>0$ for which every digraph $D$ satisfies $\vec{\chi}(D)\leq\lceil(1-\varepsilon)(\tilde{\Delta}(D)+1)+\varepsilon\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)\rceil$. This implies both Reed’s result and an independent result of Harutyunyan and Mohar for oriented graphs.
To obtain this upper bound on $\vec{\chi}$, we prove that every digraph $D$ with ${\omega}(D)>\frac{2}{3}(\Delta_{\max}(D)+1)$, where $\Delta_{\max}(D)=\max_{v\in V(D)}\max(d^{+}(v),d^{-}(v))$, admits an acyclic set of vertices intersecting each biclique of $D$, which generalises a result of King.
Keywords: Dichromatic number, dicolouring, biclique number.
@@ 1 Introduction
Conjecture 1.1 (Reed [31]).
Every graph $G$ satisfies $\chi(G)\leq\left\lceil\frac{\Delta(G)+1+\omega(G)}{2}\right\rceil$.
@@
$\chi(G)\leq\lceil(1-\varepsilon)(\Delta(G)+1)+\varepsilon\omega(G)\rceil.$
Observe that Conjecture 1.1 is exactly Theorem 1.2 for $\varepsilon=\frac{1}{2}$. Restricted to graphs with sufficiently large maximum degree, Hurley et al. [21] proved that Theorem 1.2 holds for $\varepsilon\leq 0.119$, improving on earlier results obtained successively by Bonamy et al. [9] and by Delcourt and Postle [12].
The purpose of this work is to propose analogues of Conjecture 1.1 and Theorem 1.2 for digraphs. We thus need three digraph parameters that extend $\Delta$, $\omega$, and $\chi$ respectively. By extension, we mean that, if $D$ is a symmetric digraph, these digraph parameters on $D$ must coincide with $\Delta$, $\omega$, and $\chi$ on the underlying graph of $D$. One can easily check that this is actually the case for $\tilde{\Delta}$, ${\omega}$, and $\vec{\chi}$, defined as follows.
Let $D$ be a digraph. We let $\tilde{\Delta}(D)$ denote the maximum geometric mean of the in- and out-degrees of the vertices of $D$, that is $\tilde{\Delta}(D)=\max\{\sqrt{d^{+}(v)\cdot d^{-}(v)}\mid v\in V(D)\}$. A biclique of $D$ is a set of vertices inducing a complete digraph. The biclique number ${\omega}(D)$ of $D$ is the size of the largest biclique of $D$. The notions of dicolouring and dichromatic number were introduced in the late 1970s by Erdős and Neumann-Lara [28, 13]. A $k$-dicolouring of $D$ is a function $c:V(D)\to[k]$ such that $c^{-1}(i)$ induces an acyclic subdigraph on $D$ for each $i\in[k]$. The dichromatic number of $D$, denoted by $\vec{\chi}(D)$, is the smallest $k\in\mathbb{N}$ such that $D$ admits a $k$-dicolouring. These two concepts were rediscovered by Mohar in the 2000s [26, 8] and received a lot of attention since then. It turned out that many classical results on graph colouring can be extended to digraphs with these notions, which shows that dicolouring is somehow the right extension of proper colouring to digraphs. For instance, people generalised the Strong Perfect Graph Theorem [6], Gyárfás-Sumner’s Conjecture [4, 3, 11], or Brooks-type results [18, 1, 2, 16].
@@
Conjecture 1.3 (Harutyunyan and Mohar [19]).
@@
$\vec{\chi}(D)\leq\left\lceil\frac{\tilde{\Delta}(D)}{2}\right\rceil+1.$
As a partial result towards their conjecture, they proved the following for $\varepsilon=e^{-13}$.
Theorem 1.4 (Harutyunyan and Mohar [19]).
There exists $\varepsilon>0$ such that every oriented graph $D$ with $\tilde{\Delta}(D)$ large enough satisfies $\vec{\chi}(D)\leq(1-\varepsilon)\tilde{\Delta}(D)$.
@@
Every digraph $D$ satisfies $\vec{\chi}(D)\leq\left\lceil\frac{\tilde{\Delta}(D)+1+\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)}{2}\right\rceil$.
@@
$\vec{\chi}(D)\leq\lceil(1-\varepsilon)(\tilde{\Delta}(D)+1)+\varepsilon\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)\rceil.$
@@
Our proof of Theorem 1.6 is inspired by a short proof of Theorem 1.2 due to King and Reed [24]. Along the way, we generalise to digraphs a collection of intermediate results that we find of independent interest. We refer the reader to the recent book of Stiebitz, Schweser, and Toft [32, Chapter 6] for a complete proof of Theorem 1.2 which includes proofs of all intermediate results (on undirected graphs) mentioned later on.
The paper is organised as follows. In Section 2, we give notation on digraphs and preliminary results used in the paper. In Section 3, using the probabilistic method, we prove the following upper bound on the dichromatic number of sparse digraphs, which generalises to digraphs a classical result of Molloy and Reed [27]. Given a digraph $D$ with maximum degree $\Delta_{\max}(D)=\Delta$ and a vertex $v$ of $D$, we denote by $m^{+}(v)$ and $m^{-}(v)$ the number of arcs of $D[N^{+}(v)]$ and $D[N^{-}(v)]$ respectively. We say that $D$ is $B$-sparse if, for every vertex $v\in V(D)$, we have $\min(m^{+}(v),m^{-}(v))\leq\Delta(\Delta-1)-B$.
@@
There exists an integer $\Delta_{0}\geq 1$ such that the following holds. For every digraph $D$ with maximum degree $\Delta_{\max}(D)=\Delta>\Delta_{0}$, if $D$ is $B$-sparse with $B>\Delta\log^{3}\Delta$, then
$\vec{\chi}(D)\leq(\Delta+1)-\frac{B}{4e^{7}\Delta}.$
@@
Let $D$ be a digraph with ${\omega}(D)>\frac{2}{3}(\Delta_{\max}(D)+1)$. Then $D$ has an acyclic set of vertices $I$ such that ${\omega}(D-I)=\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)-1$.
This implies that, in a minimum counterexample of Theorem 1.6, there must be a significant gap between the biclique number and the maximum degree. We indeed prove a more general result, as we characterise exactly the digraphs $D$ with ${\omega}(D)\geq\frac{2}{3}(\Delta_{\max}(D)+1)$ that do not admit such an acyclic set of vertices. We thus extend a result due to Christofides et al. [10]. We also discuss the consequences of our result on list-dicolouring of digraphs, and show that we generalise a result due to Haxell [20] that we find of independent interest.
In Section 5, we deal with vertices having a dense neighbourhood in a counterexample of Theorem 1.6. We extend a result of King and Reed [24] by proving the following.
@@
Let $a$ and $\varepsilon$ be constants satisfying $0<\varepsilon\leq\frac{1}{6}-4\sqrt{a}$. There exists $\Delta(a)\geq 1$ such that the following holds. Let $D$ be a digraph with maximum degree $\Delta_{\max}(D)=\Delta\geq\Delta(a)$ satisfying ${\omega}(D)\leq\frac{2}{3}(\Delta+1)$ and let $v\in V(D)$ be such that $\max(m^{-}(v),m^{+}(v))>(1-a)\Delta(\Delta-1)$. Then
$\vec{\chi}(D)\leq\max\left(\vec{\chi}(D-v),(1-\varepsilon)(\Delta+1)\right).$
@@ 2.1 General notation on digraphs
Let $D$ be a digraph. The underlying graph of $D$, denoted by $\operatorname{UG}(D)$, is the undirected graph with vertex-set $V(D)$ in which $uv$ is an edge if and only if $uv$ or $vu$ is an arc of $D$. The symmetric part of $D$, denoted by $S(D)$, is the undirected graph with vertex-set $V(D)$ in which $uv$ is an edge if and only if $\{uv,vu\}$ is a digon of $D$. A matching of $D$ is a set of pairwise disjoint arcs of $D$.
@@
The maximum geometric mean of $D$ is defined as $\tilde{\Delta}(D)=\max_{v\in V(D)}\sqrt{d^{-}(v)\cdot d^{+}(v)}$. By definition, observe that we have $\Delta_{\min}(D)\leq\tilde{\Delta}(D)\leq\Delta_{\max}(D)$. We say that $D$ is $\Delta$-diregular if, for every vertex $v\in V(D)$, $d^{-}(v)=d^{+}(v)=\Delta$. Note that $D$ being diregular implies $\Delta_{\min}(D)=\Delta_{\max}(D)$ but the converse is not true.
Let $X,Y$ be two disjoint set of vertices of $D$. We denote by $A(X,Y)$ the set of arcs $uv\in A(D)$ such that $u\in X$ and $v\in Y$.
A list assignment $L$ of $D$ is a function that associates a list of colours to every vertex $v$ of $D$. It is a $k$-list assignment if $|L(v)|\geq k$ holds for every vertex $v$ of $D$. Let $L$ be a list assignment. An $L$-dicolouring of $D$ is a dicolouring $\alpha$ of $D$ such that, for every vertex $v$ of $D$, $\alpha(v)\in L(v)$. We say that $D$ is $L$-dicolourable if $D$ admits an $L$-dicolouring. Given an integer $k$, we say that $D$ is $k$-dichoosable if, for every $k$-list assignment $L$ of $D$, $D$ is $L$-dicolourable.
@@ 2.2 Probabilistic tools
Let $A_{1},A_{2},\dots,A_{n}$ be events in an arbitrary probability space. Suppose that each event $A_{i}$ is mutually independent of a set of all the other events but at most $d$, and that $\mathbf{Pr}(A_{i})\leq p$ for all $1\leq i\leq n$. If
$ep(d+1)\leq 1,$
then $\mathbf{Pr}\left(\bigwedge_{i=1}^{n}\overline{A_{i}}\right)>0$.
The second result we need is the celebrated concentration bound due to Talagrand [33]. We use it in the following form (see [22, Theorem 6.3]).
@@
Let $X$ be a non-negative random variable, not identically 00, which is determined by $n$ independent trials and satisfying the following for some $c,r>0$:
@@
    for any $s$, if $X\geq s$ then there is a set of at most $rs$ trials whose outcomes certify that $X\geq s$.
@@
$\mathbf{Pr}\left(|X-\mathbf{E}(X)|>t\right)\leq 4\exp\left(-\frac{t^{2}}{8rc^{2}(4\mathbf{E}(X)+t)}\right)$
for any real number $t$ with $t>96c\sqrt{r\mathbf{E}(X)}+128rc^{2}$.

### v3

@@ Abstract
In 1998, Reed conjectured that every graph $G$ satisfies $\chi(G)\leqslant\lceil\frac{1}{2}(\Delta(G)+1+\omega(G))\rceil$. As a partial result, he proved the existence of $\varepsilon>0$ such that every graph $G$ satisfies $\chi(G)\leqslant\lceil(1-\varepsilon)(\Delta(G)+1)+\varepsilon\omega(G)\rceil$. We propose an analogous conjecture for digraphs. Given a digraph $D$, we denote by $\vec{\chi}(D)$ the dichromatic number of $D$, which is the minimum number of colours needed to partition $D$ into acyclic induced subdigraphs. We let $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)$ denote the order of a largest biclique (a set of vertices inducing a complete digraph) of $D$ and $\tilde{\Delta}(D)=\max_{v\in V(D)}\sqrt{d^{+}(v)\cdot d^{-}(v)}$. We conjecture that every digraph $D$ satisfies $\vec{\chi}(D)\leqslant\lceil\frac{1}{2}(\tilde{\Delta}(D)+1+\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D))\rceil$, which if true implies Reed’s conjecture. As a partial result, we prove the existence of $\varepsilon>0$ for which every digraph $D$ satisfies $\vec{\chi}(D)\leqslant\lceil(1-\varepsilon)(\tilde{\Delta}(D)+1)+\varepsilon\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)\rceil$. This implies both Reed’s result and an independent result of Harutyunyan and Mohar for oriented graphs.
To obtain this upper bound on $\vec{\chi}$, we prove that every digraph $D$ whose biclique number is larger than $\frac{2}{3}(\Delta_{\max}(D)+1)$ admits an acyclic set of vertices intersecting each biclique of $D$, where $\Delta_{\max}(D)=\max_{v\in V(D)}\max(d^{+}(v),d^{-}(v))$. This generalises a result of King.
We finally give a short proof that all oriented graphs $D$ satisfy $\vec{\chi}(D)\leqslant\frac{\sqrt{2}}{2}\tilde{\Delta}(D)+2$, improving on a result of Golowich.
Keywords: Dichromatic number, maximum degree, biclique number.
@@ 1 Introduction
Conjecture 1.1 (Reed [34, Conjecture 1]).
Every graph $G$ satisfies $\chi(G)\leqslant\left\lceil\frac{\Delta(G)+1+\omega(G)}{2}\right\rceil$.
@@
$\chi(G)\leqslant\lceil(1-\varepsilon)(\Delta(G)+1)+\varepsilon\omega(G)\rceil.$
Observe that Conjecture 1.1 is exactly Theorem 1.2 for $\varepsilon=\frac{1}{2}$. Restricted to graphs with sufficiently large maximum degree, Hurley et al. [23] proved that Theorem 1.2 holds for $\varepsilon\leqslant 0.119$, improving on earlier results obtained successively by Bonamy et al. [9] and by Delcourt and Postle [12].
The purpose of this work is to propose analogues of Conjecture 1.1 and Theorem 1.2 for digraphs. We thus need three digraph parameters that extend $\Delta$, $\omega$, and $\chi$ respectively. By extension, we mean that, if $D$ is a symmetric digraph, these digraph parameters on $D$ must coincide with $\Delta$, $\omega$, and $\chi$ on the underlying graph of $D$. One can easily check that this is actually the case for $\tilde{\Delta}$, $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}$, and $\vec{\chi}$, defined as follows.
Let $D$ be a digraph. We let $\tilde{\Delta}(D)$ denote the maximum geometric mean of the in- and out-degrees of the vertices of $D$, that is $\tilde{\Delta}(D)=\max\{\sqrt{d^{+}(v)\cdot d^{-}(v)}\mid v\in V(D)\}$. A biclique of $D$ is a set of vertices inducing a complete digraph. The biclique number $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)$ of $D$ is the order of the largest biclique of $D$. The notions of dicolouring and dichromatic number were introduced in the late 1970s by Erdős and Neumann-Lara [31, 13]. A $k$-dicolouring of $D$ is a function $c:V(D)\to[k]$ such that $c^{-1}(i)$ induces an acyclic subdigraph on $D$ for each $i\in[k]$. The dichromatic number of $D$, denoted by $\vec{\chi}(D)$, is the smallest $k\in\mathbb{N}$ such that $D$ admits a $k$-dicolouring. These two concepts were rediscovered by Mohar in the 2000s [29, 8] and received a lot of attention since then. It turned out that many classical results on graph colouring can be extended to digraphs with these notions, which shows that dicolouring is somehow the right extension of proper colouring to digraphs. For instance, people generalised the Strong Perfect Graph Theorem [6], Gyárfás-Sumner’s Conjecture [4, 11, 3], and Brooks-type results [20, 1, 2, 17].
@@
Conjecture 1.3 (Harutyunyan and Mohar [21, Conjecture 1.5]).
@@
$\vec{\chi}(D)\leqslant\left\lceil\frac{\tilde{\Delta}(D)}{2}\right\rceil+1.$
As a partial result towards their conjecture, they obtained the following.
Theorem 1.4 (Harutyunyan and Mohar [21, Theorem 2.1]).
Every oriented graph $D$ with $\tilde{\Delta}(D)$ large enough satisfies $\vec{\chi}(D)\leqslant(1-e^{-13})\tilde{\Delta}(D)$.
Theorem 1.4 have been strengthened by Golowich [16], who obtained that all oriented graphs $D$ satisfy $\vec{\chi}(D)\leqslant\sqrt{2/3}\tilde{\Delta}(D)+O(1)$. Combining Golowich’s idea with a recent result of the second author, in Section 3 we give a short proof of the following stronger bound, thereby approaching Conjecture 1.3.
Theorem 1.5.
Let $D$ be an oriented graph with underlying graph $G$. Then $\vec{\chi}(D)\leqslant\frac{1}{3}\Delta(G)+2$, $\vec{\chi}(D)\leqslant\frac{2}{3}\Delta_{\max}(D)+2$, and $\vec{\chi}(D)\leqslant\frac{\sqrt{2}}{2}\tilde{\Delta}(D)+2$.
@@
Every digraph $D$ satisfies $\vec{\chi}(D)\leqslant\left\lceil\frac{\tilde{\Delta}(D)+1+\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)}{2}\right\rceil$.
@@
$\vec{\chi}(D)\leqslant\lceil(1-\varepsilon)(\tilde{\Delta}(D)+1)+\varepsilon\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)\rceil.$
@@
Our proof of Theorem 1.7 is inspired by a short proof of Theorem 1.2 due to King and Reed [26]. Along the way, we generalise a collection of intermediate results to digraphs that we find of independent interest. We refer the reader to the recent book of Stiebitz, Schweser, and Toft [35, Chapter 6] for a complete proof of Theorem 1.2 which includes proofs of all intermediate results (on undirected graphs) mentioned later on.
In Section 4, using the probabilistic method, we prove the following upper bound on the dichromatic number of sparse digraphs, which generalises to digraphs a classical result of Molloy and Reed [30, Theorem 10.5]. Given a digraph $D$ with maximum degree $\Delta_{\max}(D)=\Delta$ and a vertex $v$ of $D$, we denote by $m^{+}(v)$ and $m^{-}(v)$ the number of arcs of $D[N^{+}(v)]$ and $D[N^{-}(v)]$ respectively. We say that $D$ is $B$-sparse if, for every vertex $v\in V(D)$, we have $\min(m^{+}(v),m^{-}(v))\leqslant\Delta(\Delta-1)-B$.
@@
There exists an integer $\Delta_{0}\geqslant 1$ such that the following holds. For every digraph $D$ with maximum degree $\Delta_{\max}(D)=\Delta>\Delta_{0}$, if $D$ is $B$-sparse with $B>\Delta\log^{3}\Delta$, then
$\vec{\chi}(D)\leqslant(\Delta+1)-\frac{B}{4e^{7}\Delta}.$
@@
Let $D$ be a digraph with $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)>\frac{2}{3}(\Delta_{\max}(D)+1)$. Then $D$ has an acyclic set of vertices $I$ such that $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D-I)=\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)-1$.
This implies that, in a minimum counterexample to Theorem 1.7, there must be a significant gap between the biclique number and the maximum degree. We indeed prove a more general result, as we characterise exactly the digraphs $D$ with $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)\geqslant\frac{2}{3}(\Delta_{\max}(D)+1)$ that do not admit such an acyclic set of vertices. We thus extend a result due to Christofides et al. [10]. We also discuss the consequences of our result on list-dicolouring of digraphs, and show a generalisation of a result due to Haxell [22] that we find of independent interest.
In Section 6, we deal with vertices having a dense neighbourhood in a counterexample to Theorem 1.7. We extend a result of King and Reed [26] by proving the following.
@@
Let $a$ and $\varepsilon$ be constants satisfying $0<\varepsilon\leqslant\frac{1}{6}-4\sqrt{a}$. There exists $\Delta(a)\geqslant 1$ such that the following holds. Let $D$ be a digraph with maximum degree $\Delta_{\max}(D)=\Delta\geqslant\Delta(a)$ satisfying $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)\leqslant\frac{2}{3}(\Delta+1)$ and let $v\in V(D)$ be such that $\max(m^{-}(v),m^{+}(v))>(1-a)\Delta(\Delta-1)$. Then
$\vec{\chi}(D)\leqslant\max\left(\vec{\chi}(D-v),(1-\varepsilon)(\Delta+1)\right).$
@@ 2.1 General notation on digraphs
Let $D$ be a digraph. The underlying graph of $D$, denoted by $\operatorname{UG}(D)$, is the undirected graph with vertex-set $V(D)$ in which $uv$ is an edge if and only if $uv$ or $vu$ is an arc of $D$. The symmetric part of $D$, denoted by $S(D)$, is the undirected graph with vertex-set $V(D)$ in which $uv$ is an edge if and only if $\{uv,vu\}$ is a digon of $D$. A matching of $D$ is a set of pairwise disjoint arcs of $D$. The complement $\bar{D}$ of $D$ is the digraph with vertex-set $V(D)$ and arc-set $(V(D)\times V(D))\setminus A(D)$.
@@
The maximum geometric mean of $D$ is defined as $\tilde{\Delta}(D)=\max_{v\in V(D)}\sqrt{d^{-}(v)\cdot d^{+}(v)}$. By definition, observe that we have $\Delta_{\min}(D)\leqslant\tilde{\Delta}(D)\leqslant\Delta_{\max}(D)$. We say that $D$ is $\Delta$-diregular if, for every vertex $v\in V(D)$, $d^{-}(v)=d^{+}(v)=\Delta$. Note that $D$ being diregular implies $\Delta_{\min}(D)=\Delta_{\max}(D)$ but the converse is not true.
Let $X,Y$ be two disjoint sets of vertices of $D$. We denote by $A(X,Y)$ the set of arcs $uv\in A(D)$ such that $u\in X$ and $v\in Y$.
A list assignment $L$ of $D$ is a function that associates a list of colours to every vertex $v$ of $D$. It is a $k$-list assignment if $|L(v)|\geqslant k$ holds for every vertex $v$ of $D$. Let $L$ be a list assignment. An $L$-dicolouring of $D$ is a dicolouring $\alpha$ of $D$ such that, for every vertex $v$ of $D$, $\alpha(v)\in L(v)$. We say that $D$ is $L$-dicolourable if $D$ admits an $L$-dicolouring. Given an integer $k$, we say that $D$ is $k$-dichoosable if, for every $k$-list assignment $L$ of $D$, $D$ is $L$-dicolourable.
@@ 2.2 Probabilistic tools
Let $A_{1},A_{2},\dots,A_{n}$ be events in an arbitrary probability space. Suppose that each event $A_{i}$ is mutually independent of a set of all the other events but at most $d$, and that $\mathbb{P}(A_{i})\leqslant p$ for all $1\leqslant i\leqslant n$. If
$ep(d+1)\leqslant 1,$
then $\mathbb{P}\left(\bigwedge_{i=1}^{n}\overline{A_{i}}\right)>0$.
The second result we need is the celebrated concentration bound due to Talagrand [36]. We use it in the following form (see the appendix of [19]).
@@
Let $X$ be a random variable valued in $\mathbb{N}$, determined by $n$ independent trials and satisfying the following for some integers $c,r\geqslant 1$:
@@
    for every $s\in\mathbb{N}$, if $X\geqslant s$ then there is a set of at most $rs$ trials whose outcomes certify that $X\geqslant s$.
@@
$\mathbb{P}\left(|X-\mathbb{E}(X)|>t\right)\leqslant 4\exp\left(\frac{-t^{2}}{32c^{2}r(\mathbb{E}(X)+t)}\right)$
for any real number $t>126c\sqrt{r\mathbb{E}(X)}+344c^{2}r$.
3 A short proof for oriented graphs
In this section we give a short proof of Theorem 1.5, that we first recall here for convenience.
See 1.5
Our proof relies on a combination of the two following known results.
Theorem 3.1 (Lovász [28, Theorem 1]).
Let $G$ be a graph and $d_{1}\geqslant\ldots\geqslant d_{p}\geqslant 1$ be a sequence of integers such that $\sum_{i=1}^{p}d_{i}\leqslant\Delta(G)-p+1$. Then $G$ admits a partition $(V_{1},\ldots,V_{p})$ of its vertex-set such that, for every $i\in[k]$, $\Delta(G[V_{i}])\leqslant d_{i}$.
Theorem 3.2 (Picasarri-Arrieta [32, Corollary 8]).
Every oriented graph $D$ satisfies
$\vec{\chi}(D)\leqslant\max(2,\Delta_{\min}(D)).$
Proof of Theorem 1.5.
Let $D$ be any oriented graph with underlying graph $G$. By Theorem 3.1 applied with $p=\lceil(\Delta+1)/6\rceil$ and $d_{1}=\ldots=d_{p}=5$, $G$ admits a partition $(V_{1},\dots,V_{p})$ of its vertex-set such that $\Delta(G[V_{i}])\leqslant 5$.
For every $i\in[p]$, we obtain that $D[V_{i}]$ is an orientation of a graph with maximum degree at most $5$. In particular, it follows that $\Delta_{\min}(D[V_{i}])\leqslant 2$. By Theorem 3.2, we thus have $\vec{\chi}(D[V_{i}])\leqslant 2$. We thus obtain
$\vec{\chi}(D)\leqslant\sum_{i=1}^{p}\vec{\chi}(D[V_{i}])\leqslant 2p\leqslant\frac{\Delta(G)}{3}+2.$ (1)
It is straightforward that $\Delta(G)\leqslant 2\Delta_{\max}(D)$, and it follows that $\vec{\chi}(D)\leqslant\frac{2}{3}\Delta_{\max}(D)+2$.
With Equation (1) in hand, we now prove the last inequality of the statement. Assume for a contradiction that there exist a real number $\tilde{\Delta}$ and a digraph $D$ with $\tilde{\Delta}(D)\leqslant\tilde{\Delta}$ and $\vec{\chi}(D)>\left\lfloor\frac{\sqrt{2}}{2}\tilde{\Delta}+2\right\rfloor=k$. Among all such digraphs $D$, we choose one with minimum order. This implies that all vertices $v\in V(D)$ satisfy
$\min(d^{-}(v),d^{+}(v))\geqslant k-1\geqslant\frac{\sqrt{2}}{2}\tilde{\Delta}$,
for otherwise a (k1)(k-1)-dicolouring of DvD-v (which exists by minimality of $D$) can be extended to $D$ by choosing for $v$ a colour that is not appearing in its in- or out-neighbourhood.
Let $v$ be any vertex for which $d^{-}(v)+d^{+}(v)$ is maximised, that is $d^{-}(v)+d^{+}(v)=\Delta(G)$, where $G$ is the underlying graph of $D$. We assume by directional duality that $d^{-}(v)\geqslant d^{+}(v)$, and in particular $d^{+}(v)\leqslant\tilde{\Delta}$. By definition of $\tilde{\Delta}$, we have $d^{-}(v)\leqslant\frac{\tilde{\Delta}^{2}}{d^{+}(v)}$.
Let $f\colon x\mapsto x+\frac{\tilde{\Delta}^{2}}{x}$. Observe that, over the interval $\frac{\sqrt{2}}{2}\tilde{\Delta},\tilde{\Delta}$], $f$ is maximised for $x=\frac{\sqrt{2}}{2}\tilde{\Delta}$. This shows
$\Delta(G)=d^{-}(v)+d^{+}(v)\leqslant d^{+}(v)+\frac{\tilde{\Delta}^{2}}{d^{+}(v)}\leqslant\frac{3\sqrt{2}}{2}\tilde{\Delta}$.
It follows from (1) and the inequality above that
$\vec{\chi}(D)\leqslant\frac{\sqrt{2}}{2}\tilde{\Delta}+2.$
Since $\vec{\chi}(D)$ is an integer, it follows that $\vec{\chi}(D)\leqslant k$, a contradiction. ∎